preferred? Assume repel 6-34. Potable water is in short supply in many countries To address this need, two mutually exclusive water purification systems are being considered for implementation in China. Doing nothing is not an option. Refer to the data below and state your key assumptions in working this problem (6.5) System 7 System 2 Capital investment $100,000 $150,000 Annual revenues 50,000 Annual expenses MV at end of useful life 20,000 Useful life 70,000 22,000 40,000 5 years 16.5% 10 yearsUs | | 15.0% IRR a. Use the PW method to determine which system should be selected When MARR 8% per year. ll. Which system should be selected when MARR-,15% per year? c.. 2
Answer #1

1) Solution: Option-2 should be selected

Working:

PW(1) = -$100,000 - $80,000 (P/F,8%,5) + $20,000 (P/F,8%,10) + $28,000 (P/A,8%,10) = $42,699

PW(2) = -$150,000 + $30,000 (P/A,8%,10) = 51,303

Thus Option-2 should be selected

2) Solution: Option-1 should be selected

Working:

PW(1) = -$100,000 - $80,000 (P/F,15%,5) + $20,000 (P/F,15%,10) + $28,000 (P/A,15%,10)

= -$100,000 - $39,774.13882 + $4943.69412 + $140,525.52

= $5695.075

PW(2) = -$150,000 + $30,000 (P/A,15%,10) = $563

Thus Option-1 should be selected